John buys stamps. Some are 33 cents and the rest are 20 cents. The total equals 4.50.
Answer by ankor@dixie-net.com(15656)

Some are 33 cents and the rest are 20 cents. The total equals 4.50.
:
Let x = no. of 33 cent stamps, y = no. 20 cent stamps, keep everything in cents
33x + 20y = 450
:
Assuming there are at least 2 ea 20 cent stamps: (450-40)/33 = 12.4
We can assume the number of 33 stamps is less than 13
Try 12: 33*12 = 396, 450 - 396 = 54, not a multiply of 20
assume it has to be an even number
Try 10: 33*10 = 330, 450 - 330 = 120, bingo!
:
We have 10 ea 33 cent and 6 ea 20 cent stamps